How many gallons of a 50% antifreeze solution must be mixed with 100 gallons of a 30% antifreeze solution to get a mixture that is 40% antifreeze?
Answer by checkley79(3341)

.50X+.30*100=.40(100+X)
.50X+30=40-.40X
.50X-.40X=40-30
.10X=10
X=10/.10
X=100 GALLONS OF 50% ARE NEEDED.
PROOF:
.50*100+.30*100=.40(100+100)
50+30=.40*200
80=80